OpenStudy (anonymous):

If you just totally help me Ill give you a medal and fan ya.. Find the slope of the line that passes through (-9, 5) and (6, 13)

OpenStudy (anonymous):

Slope = change of y/change of x Look at the y-values: the change in y-values is (13-5) = 8 Change in x-values is (6 - (-9)) = 15 slope = 8/15